# Deep learning and whole-brain networks for biomarker discovery: modeling the dynamics of brain fluctuations in resting-state and cognitive tasks

### v1

@@ Abstract
Results: Bifurcation parameter distributions differed significantly across task and resting-state conditions ($p<0.0001$ for all but one comparison). Task-based brain states exhibited higher bifurcation values compared to rest.
@@ 1 Introduction
Non-invasive neuroimaging techniques are becoming widespread, providing a rich information source that revolutionized the study of brain dynamics, enabling researchers to correlate brain activity with observable behaviors and cognitive processes among other goals [1, 2]. Functional magnetic resonance imaging (fMRI), in particular, provides detailed high-resolution spatiotemporal data of oxygen consumption, arguably related to regional brain activity. However, interpreting these data remains a major challenge, as the complexity of brain signals often requires advanced computational approaches and an adequate understanding thereof to uncover meaningful patterns or to elucidate among possible interpretations. Deep learning has emerged as a powerful tool for this purpose [3, 4], offering flexible models that can learn complex structures directly from raw data without extensive handcrafted features –they learn directly from the available training samples. These models excel at extracting intricate patterns across diverse domains and have demonstrated their versatility in a wide range of tasks, including classifying handwritten digits [5], detecting objects in a scene [6], segmenting brain tumors [7], playing board games [8] and bidirectional mapping of structural and functional brain connectivity [9].
@@
Biomarkers play a crucial role in advancing 4P medicine (predictive, preventive, personalized, and participatory), enabling, among other things, an early detection of health-compromising conditions (especially those preventable or treatable at an early stage), monitoring disease progression, evaluating the effectiveness of treatments, and tailoring treatments to individual needs. Given that 4P medicine is inherently participative, it necessitates the collection of extensive and diverse data from a broad range of individuals to ensure comprehensive insights and effective implementation of personalized healthcare strategies. For this reason, there is an increasing need in neuroscience for adequate and useful biomarkers [10], defined as measurable biological indicators that provide insights to distinguish normal from pathological processes in the brain, or responses to interventions such as therapies or drugs. For instance, fMRI-based biomarkers can detect functional connectivity abnormalities distinguishing Alzheimer’s disease from mild cognitive impairment and healthy controls [11]. Similarly, Bruin et al. [12] proposed biomarkers to predict outcomes of electroconvulsive therapy, enabling more precise interventions for patients with treatment-resistant depression.
@@
In this context, whole-brain computational models represent a promising approach to understanding the dynamic processes of the brain. These models simulate large-scale brain networks, capturing complex neural interactions throughout the brain. Whole-brain models have been valuable in various applications, including understanding network-level disruptions in neurological disorders [13], guiding potential therapeutic interventions [14], and optimizing simulations to achieve high fidelity with limited computational resources [15]. An important example is the study by Deco et al. [16] where the parameters of a simple brain network model were optimized, identifying a dynamical cortical core responsible for driving activity across the brain. Although this approach provided valuable insights, the model parameters were not considered as biomarkers for specific brain states or cognitive functions, thereby limiting their generalizability and broader applicability.
@@
To this avail, the present study aims to bridge this gap using a whole-brain computational model to explore whether constructive parameters can be biomarkers for different task-based brain states. In particular, we focus on a neural mass model, a mathematical framework that simulates the activity of large groups of neurons, which is capable of capturing the fluctuations in brain activity and transitions between brain states [17, 18]. This model enables to characterize the brain’s dynamical system through bifurcation parameters, which describe the shifts in the brain’s stability and oscillatory patterns. Our claim is that these bifurcation parameters may be sufficient to characterize distinct task-based brain states associated with various cognitive tasks. We trained deep learning models to predict bifurcation parameters using synthetic BOLD signals generated with Hopf’s brain network model (see Fig. 1). By leveraging synthetic data, we overcome the limitations of scarce labeled datasets and ensure that our models are trained on a standardized, well-controlled set of inputs. We then applied the trained models to the Human Connectome Project (HCP) dataset to estimate bifurcation parameter distributions across cohorts performing different cognitive tasks. By comparing these distributions, we investigate whether bifurcation parameters can serve as meaningful brain state indicators, potentially offering a novel approach to mapping cognitive functions through model-driven biomarkers.
@@ 2 Results
This study examined brain dynamics across multiple cognitive tasks using data from the HCP, covering seven task-based conditions and resting-state data. Brain imaging data were processed using standard pipelines and segmented into 80 regions of interest (ROIs) with cortical and subcortical coverage using the DBS80 parcellation. We used a brain network model based on the supercritical Hopf bifurcation to model neural activity, simulating interactions among these brain regions. Model parameters, such as global coupling factor ($G$) and intrinsic oscillation frequencies ($\omega_{j}$), were calibrated to match the phase functional connectivity dynamics (FCD) observed in the HCP resting-state data. Given the limited sample size for training a deep learning model, synthetic BOLD signals were generated to create a voluminous dataset, allowing these models to be effectively trained. Thus, deep learning models were trained on synthetic data to predict the model bifurcation parameters $a_{j}$ for each brain region $j$ from BOLD signals. These values serve as a proxy for brain state characteristics, capturing task-related variations in neural activity across brain regions. The trained models were then applied to empirical HCP data, allowing statistical analysis to identify significant differences across tasks and brain networks, thus providing insights into the underlying brain states associated with each cognitive task.
Figure 2: Determination of optimal coupling factor $G$ and statistical comparison of bifurcation parameters across cohorts. (A) The optimal global coupling factor $G$ was estimated through a parameter sweep, minimizing the Kolmogorov-Smirnov distance between the FCD of empirical and simulated BOLD data. The polynomial fit identifies the $G$ value that best aligns the simulated data with the observed resting-state FCD patterns. (B) Distribution of mean bifurcation parameter values across different cognitive task and resting-state cohorts. Statistical significance between cohorts is assessed using the Mann-Whitney-Wilcoxon test, with Benjamini-Hochberg correction applied to control for multiple comparisons. Statistical significance is indicated by **** for $p<=0.0001$, *** for $0.0010.0001<p<=$ 0.001, ** for $0.010.001<p<=$ 0.01, * for $0.050.01<p<=$ 0.05 and "ns" for $10.05<p<=$ 1. These results reveal distinguishable brain state characteristics associated with each cognitive task.
2.1 Developing of the deep learning models
@@ 2.1.2 Samples are more important than time steps
We conducted a comprehensive hyperparameter exploration to understand how the dimensions of the synthetic BOLD signals impact the predictive performance of our deep learning models. In this exploration, the number of samples ($S$), nodes ($N$), and time steps ($W$) were jointly adjusted, while ensuring the total size of the dataset did not exceed 160 million values ($S\times N\times W$) to maintain a limited computational budget. Since the selected parcellation fixes the number of nodes at $N=80$, we focused on just varying the values of $S$ and $W$. For each viable combination of these parameters (i.e., those that remained within the imposed threshold), a separate model was trained using the time series approach. The performance metrics of these models are reported in Table 1.
Our findings confirm that increasing the number of samples consistently improves model performance across all tested configurations. This was expected, as more samples generally provide better results when training deep learning models [19]. However, increasing the number of time steps shows inconsistent improvements. Although in some cases a slight boost in performance is noticeable with higher $W$ values, the effect plateaus quickly, and further $W$ increments indeed diminish the performance. These results suggest that the most computationally efficient strategy maximizes the number of samples while keeping the time steps relatively low. Importantly, we observed a performance degradation when the time step count dropped to $W=25$, indicating that reducing $W$ to this point negatively affects the model’s accuracy. Based on these insights, we selected $W=50$ and $S=40,000$ as the optimal working point for all subsequent analyses, balancing performance and computational efficiency.
@@ 2.2 Analysis of predicted bifurcation parameters from the HCP dataset
2.2.1 Cohort separability
Inference on the HCP dataset reveals distinct patterns in the distribution of mean bifurcation parameters across cohorts, supporting their separability based on brain state characteristics. As detailed in the Methods section, we conducted pairwise comparisons of these distributions, with the results presented in Fig. 2B as violin plots with corresponding Mann-Whitney-Wilcoxon test statistics. Every cohort comparison yielded statistically significant distribution differences. Remarkably, all pairwise comparisons except one (memory vs. relational) showed a p-value below 0.0001. The plot further illustrates that most bifurcation parameter values are small negative numbers between -0.1 and 0, indicating a system preference for values in the non-oscillatory behavior of the supercritical Hopf bifurcation [16]. The resting-state cohort exhibits noticeably lower bifurcation parameter values overall, supporting the hypothesis that these parameters correlate with brain activity. This suggests that greater oscillatory dynamics are required to couple nodes and enhance inter-regional brain communication as the cognitive demands of a task increase.
@@ 2.2.2 Brain network activations vary across tasks
To explore how different brain networks are engaged across functions, we grouped cortical nodes according to their respective brain networks, as defined by the Yeo atlas, and averaged bifurcation parameter values within each cohort. After subtracting the resting-state values from each task, we obtained the difference of mean bifurcation parameters shown in Fig. 3. Overall, most bifurcation parameter values are positive, confirming that these parameters are generally higher during task engagement than at rest, and indicating a system preference for the oscillatory behavior during tasks. Also, this again supports the hypothesis that bifurcation values correlate with brain activity levels.
Interestingly, the dorsal attention network exhibits a highly variable pattern, showing the highest bifurcation values in four tasks but the lowest in two others. This variability could stem from the fact that few nodes are associated with this network, making it difficult to draw robust conclusions without further data. Excluding dorsal attention, the visual network shows the highest bifurcation values overall, ranking in the top four activations for six out of seven tasks and the top two for four tasks. This result is expected, as visual stimuli were integral to most task designs in the experiments [20]. Similarly, the default network ranks within the top four activations in all but one task, indicating its consistent engagement across different cognitive demands. In contrast, the ventral attention network shows the lowest activity, ranking last in four tasks.
When analyzing individual tasks, we found that emotion and language tasks show near-zero differences from resting-state bifurcation values, suggesting that these tasks may not be particularly demanding regarding brain activation. In contrast, the gambling, motor, and social tasks exhibit consistently positive bifurcation values across networks, suggesting higher engagement. Additionally, the relational task shows substantial variability in bifurcation values across networks, indicating diverse activation patterns.
Figure 3: Task-based differences in mean bifurcation parameters across brain networks. The mean bifurcation parameter values were calculated across cortical nodes grouped by brain network using the Yeo atlas for each task. These values were then normalized by subtracting the corresponding mean bifurcation parameter for the resting-state cohort. Each point represents the average difference per network, with colors indicating specific brain networks.
@@ 3 Discussion
We developed a novel deep-learning-based pipeline that addresses the inverse problem for an ordinary differential equation system, specifically tailored to model BOLD signals. This approach allows us to infer key dynamical parameters of a computational brain network model with remarkable efficiency and precision, relying solely on synthetic data for model training. Our results show that this model-driven inference pipeline can distinguish between different cognitive and resting states in empirical data, providing insights into the underlying brain states and supporting the use of bifurcation parameters as potential biomarkers for neural activity patterns. This capability opens a promising landscape for decoding complex task-related brain states and advancing our understanding of the human neurocognition in a computationally efficient and scalable manner.
@@
Our study addresses a significant challenge in neuroscience: large volumes of training data are typically required for deep learning models to generalize effectively [21]. By generating synthetic BOLD signals derived from a simple whole-brain computational model, we circumvented the limitations associated with scarce or heterogeneous neuroimaging data, offering a controlled, high-volume dataset for training. This approach leverages synthetic data to capture essential aspects of BOLD dynamics without reliance on extensive empirical datasets, thereby avoiding potential biases from real-world data collection conditions. This methodology strengthens the potential of our model to generalize across datasets, highlighting synthetic data as a viable solution in neuroimaging applications where data scarcity often hinders model development.
@@
Previous work in computational brain network modeling has largely focused on examining the dynamics of the human brain in resting states [22, 23, 24, 25]. In a foundational study, Deco et al. [16] demonstrated that the brain operates at maximum metastability during rest, characterized by continual transitions between functional network states. This metastable behavior is thought to enable optimal integration and segregation of information across regions, facilitating flexible cognitive processing. Building on this foundation, our work extends the application of the computational brain network model to task-related brain states, revealing that constructive parameters of the Hopf model –particularly the bifurcation parameters– can serve as biomarkers to differentiate distinct mental states induced by various cognitive tasks. Fig. 2B illustrates the increasing demand for inter-regional coordination and dynamic coupling as tasks deviate from the resting-state condition, accentuating the brain’s adaptive responses to varying levels of complexity. Additionally, the bifurcation parameters we obtained are predominantly small and negative, supporting the authors’ claim that the brain operates near the edge of bifurcation during rest.
@@
To the best of our knowledge, this study is among the earliest to adopt an image-based approach for processing BOLD time series using deep learning. To date, only Kancharala et al. has employed a similar strategy, using Gramian Angular Fields and Markov Transition Fields to encode BOLD signals into image representations [26]. That work achieved an 18% improvement in classification accuracy when distinguishing brain states induced by viewing images from standard computer vision datasets, compared to conventional time-series-based approaches. Analogously, we observed nearly a 3% decrease in normalized RMSE for bifurcation parameter prediction when using image-based transformations rather than raw time series, underscoring the advantage of this approach. By transforming temporal dynamics into spatial representations, image-based methods allow models to leverage powerful computer vision architectures adept at extracting intricate spatial features to capture complex temporal patterns across brain regions. Image-based time series methodologies in neuroimaging could pave the way for more robust and accurate tools in functional brain mapping and cognitive state classification.
@@
When analyzing the network-specific task activations shown in Fig. 3, our approach successfully links task-based activation patterns to brain networks commonly associated with these functions, aligning with previous findings in the literature [27]:
  • The default mode network is most active during introspective tasks or passive thought [28]. Since humans almost always subconsciously process the world around them, it is considered one of the most constantly active main brain networks. Consistently, we observed this network as the second most active across tasks, underscoring its ubiquitous role in internal cognitive processing.
  • The frontoparietal network (also named central executive network) enables complex problem-solving by integrating inputs from other networks for high-level cognitive functions, such as flexibility, working memory, and inhibition [29]. Our results showed that this network showed maximum activation during the memory task among all networks.

### v2

@@ Abstract
Results: Bifurcation parameter distributions differed significantly across task and resting-state conditions ($p<0.0001$ for all but two comparisons). Task-based brain states exhibited higher bifurcation values compared to rest. At the individual level, a machine learning model was able to classify the predicted bifurcation values into eight cohorts with 62.63% accuracy (well above the 12.50% chance level).
@@ 1 Introduction
Non-invasive neuroimaging techniques are becoming widespread, providing a rich information source that revolutionized the study of brain dynamics, enabling researchers to correlate brain activity with observable behaviors and cognitive processes among other goals [Fincham2020, Yang2021]. Functional magnetic resonance imaging (fMRI), in particular, provides data with high spatial resolution detailing oxygen consumption, which is related to regional brain activity. However, interpreting these data remains a major challenge, as the complexity of brain signals often requires advanced computational approaches and an adequate understanding thereof to uncover meaningful patterns or to elucidate among possible interpretations. Deep learning has emerged as a powerful tool for this purpose [Zhao2018, Thomas2023], offering flexible models that can learn complex structures directly from raw data without extensive handcrafted features —they learn directly from the available training samples. These models excel at extracting intricate patterns across diverse domains and have demonstrated their versatility in a wide range of tasks, including classifying handwritten digits [Lecun1998], detecting objects in a scene [Girshick2015], segmenting brain tumors [Havaei2017], playing board games [Silver2016], and bidirectional mapping of structural and functional brain connectivity [Jamison2024].
@@
Biomarkers play a crucial role in advancing 4P medicine (predictive, preventive, personalized, and participatory), enabling, among other things, an early detection of health-compromising conditions (especially those preventable or treatable at an early stage), monitoring disease progression, evaluating the effectiveness of treatments, and tailoring treatments to individual needs. Given that 4P medicine is inherently participative, it necessitates the collection of extensive and diverse data from a broad range of individuals to ensure comprehensive insights and effective implementation of personalized healthcare strategies. For this reason, there is an increasing need in neuroscience for adequate and useful biomarkers [Puig2020], defined as measurable biological indicators that provide insights to distinguish normal from pathological processes in the brain, or responses to interventions such as therapies or drugs. For instance, fMRI-based biomarkers can detect functional connectivity abnormalities, distinguishing Alzheimer’s disease from mild cognitive impairment and healthy controls [Khatri2022]. Similarly, Bruin et al. [Bruin2024] proposed biomarkers to predict outcomes of electroconvulsive therapy, enabling more precise interventions for patients with treatment-resistant depression.
@@
In this context, whole-brain computational models represent a promising approach to understanding the dynamic processes of the brain. These models simulate large-scale brain networks, capturing complex neural interactions throughout the brain. Whole-brain models have been valuable in various applications, including understanding network-level disruptions in neurological disorders [Pathak2022], guiding potential therapeutic interventions [Deco2014], and optimizing simulations to achieve high fidelity with limited computational resources [Herzog2024]. An important example is the study by Deco et al. [Deco2017] where the parameters of a simple brain network model were optimized, identifying a dynamical cortical core responsible for driving activity across the brain. Building on this, the field is moving towards more complex models with high-dimensional, region-specific parameters. Such approaches have been shown to significantly improve the fit to empirical data and enable the prediction of individual traits, though they require dedicated and computationally intensive optimization strategies for each subject [Wischnewski2025]. While these approaches provided valuable insights, a key unresolved question is whether these model-derived parameters could serve as biomarkers for specific brain states or cognitive functions, thereby limiting their generalizability and broader applicability.
@@
Note a key distinction: empirical data is used for selecting the optimal coupling factor and for inference, while synthetic data is used only for model training.
To this end, the present study aims to bridge this gap using a whole-brain computational model to explore whether constructive parameters can be biomarkers for different task-based brain states. In particular, we focus on a neural mass model, a mathematical framework that simulates the activity of large groups of neurons, which is capable of capturing fluctuations in brain activity and transitions between brain states [Freyer2011, Freyer2012]. This model enables the characterization of the brain’s dynamical system through bifurcation parameters, which describe the shifts in the brain’s stability and oscillatory patterns. Our claim is that these bifurcation parameters may be sufficient to characterize distinct task-based brain states associated with various cognitive tasks. We trained deep learning models to predict bifurcation parameters using synthetic BOLD signals generated with Hopf’s brain network model (see Fig. 1). By leveraging synthetic data, we overcome the limitations of scarce labeled datasets and ensure that our models are trained on a standardized, well-controlled set of inputs. We then applied the trained models to the Human Connectome Project (HCP) dataset to estimate bifurcation parameter distributions across cohorts performing different cognitive tasks. By comparing these distributions, we investigate whether bifurcation parameters can serve as meaningful brain state indicators, potentially offering a novel approach to mapping cognitive functions through model-driven biomarkers.
@@ 2 Results
This study examines brain dynamics across multiple cognitive tasks using data from the HCP, covering resting-state and seven task-based conditions. Our methodology involved a three-stage process that separated brain network model calibration, deep learning model training, and inference. The goal was to predict bifurcation parameters from BOLD signals, which serve as a proxy for brain state characteristics by capturing task-related variations in neural activity across brain regions.
First, for model calibration, we used a brain network model based on the supercritical Hopf bifurcation to simulate interactions among 80 cortical and subcortical brain regions, defined by the DK80 parcellation. The global coupling factor ($G$) of the brain network model was calibrated by matching the phase functional connectivity dynamics (FCD) of the simulated data with those observed in the empirical HCP resting-state data (preprocessed using the standard HCP pipeline [Glasser2013, Smith2013]). This ensured that our generative brain network model operates in a dynamically realistic regime.
Second, for model training, we overcome the issue of having a limited sample size of empirical data by using the calibrated brain network model to generate a voluminous synthetic BOLD dataset with known ground-truth bifurcation parameters ($a_{j}$). Deep learning models were then trained and validated exclusively on these synthetic data, using an 80/20 split.
Finally, in the inference stage, the fully trained model was applied to the empirical HCP data from all subjects and cohorts. This allowed us to perform both a group-level statistical analysis and an individual-level classification with the inferred bifurcation parameters to identify significant differences across tasks and brain networks, thus providing insights into the underlying causal dynamics [Deco2017] associated with each cognitive task.
Figure 2: Determination of optimal coupling factor $G$ and statistical comparison of bifurcation parameters across cohorts. (A) The optimal global coupling factor $G$ was estimated through a parameter sweep, minimizing the Kolmogorov-Smirnov distance between the FCD of empirical and simulated BOLD data. The polynomial fit identifies the $G$ value that best aligns the simulated data with the observed resting-state FCD patterns. (B) Distribution of mean bifurcation parameter values across different cognitive task and resting-state cohorts. Statistical significance between cohorts is assessed using the Mann-Whitney-Wilcoxon test, with Benjamini-Hochberg correction applied to control for multiple comparisons. Statistical significance is indicated by **** for $p<=0.0001$, *** for $0.0001<p<=0.001$, ** for $0.001<p<=0.01$, * for $0.01<p<=0.05$ and "ns" for $0.05<p<=1$. These results reveal distinguishable brain state characteristics associated with each cognitive task.
2.1 Development of the deep learning models
@@ 2.1.2 Samples are more important than time steps
We conducted a comprehensive hyperparameter exploration to understand how the dimensions of the synthetic BOLD signals impact the predictive performance of our deep learning models. In this exploration, the number of samples ($S$), nodes ($N$), and time steps ($W$) were jointly adjusted, while ensuring that the total size of the dataset did not exceed 160 million values ($S\times N\times W$) to maintain a limited computational budget. Since the selected parcellation fixes the number of nodes at $N=80$, we focused on just varying the values of $S$ and $W$. For each viable combination of these parameters (i.e., those that remained within the imposed threshold), a separate model was trained using the time series approach. The performance metrics of these models are reported in Table 1.
Our findings confirm that increasing the number of samples consistently improves model performance across all tested configurations. This was expected, as more samples generally provide better results when training deep learning models [Hoffmann2022]. However, increasing the number of time steps shows inconsistent improvements. Although in some cases a slight boost in performance is noticeable with higher $W$ values, the effect plateaus quickly, and further $W$ increments indeed diminish the performance. These results suggest that the most computationally efficient strategy maximizes the number of samples while keeping the time steps relatively low. Importantly, we observed a performance degradation when the time step count dropped to $W=25$, indicating that reducing $W$ to this point negatively affects the model’s accuracy. Based on these insights, we selected $W=50$ and $S=40,000$ as the optimal working point for all subsequent analyses, balancing performance and computational efficiency.
@@ 2.1.3 The image approach yields better results
2.1.4 Node ordering and parcellation choice impact model performance
To formally test whether the anatomical ordering of brain regions provided a performance advantage, we compared the performance of our original model against a null distribution. This null distribution was generated by training the same model ten separate times, each on a unique, randomly permuted ordering of the input nodes. The models trained on shuffled data achieved a normalized RMSE of 7.12 ± 0.11, with values ranging from 6.97 to 7.35. We can compare this to our original metric of 6.93, which was superior to all ten models trained on permuted data. To assess statistical significance, we conducted a one-tailed permutation test, which resulted in a p-value of 0.09, the lowest possible value achievable with ten permutations. To quantify the magnitude of this difference, we also calculated the effect size, which was found to be very large (Cohen’s d was 1.76) [Sullivan2012]. While the p-value is considered marginally significant due to the limited number of permutations, the large effect size provides strong evidence that the anatomical ordering offers a meaningful, albeit small, performance advantage. This indicates that the convolutional network is likely leveraging subtle spatial features in the anatomically ordered data, though its strong performance is not solely dependent on this specific ordering.
In addition, to assess consistency across different brain atlases, we performed a training run using the Schaefer100 parcellation. This resulted in a normalized RMSE of 8.53. It is important to interpret this result with caution, as the increasing number of nodes likely requires a longer training process to achieve optimal performance. However, this result confirms that the choice of parcellation has a tangible impact on the performance of the model and highlights the importance of atlas selection in this framework.
@@ 2.2 Analysis of predicted bifurcation parameters from the HCP dataset
2.2.1 Group-level statistical analysis
Inference on the HCP dataset reveals distinct patterns in the distribution of mean bifurcation parameters across cohorts, supporting their separability based on brain state characteristics. As detailed in the Methods section, we conducted pairwise comparisons of these distributions, with the results presented in Fig. 2B as violin plots with corresponding Mann-Whitney-Wilcoxon test statistics. Remarkably, all pairwise comparisons except one (Motor vs. Gambling) showed statistically significant distribution differences with p-values below 0.001. The plot further illustrates that most bifurcation parameter values are small negative numbers between -0.25 and 0, indicating a system preference for values in the non-oscillatory behavior of the supercritical Hopf bifurcation [Deco2017]. The resting-state cohort exhibits noticeably lower bifurcation parameter values overall, supporting the hypothesis that these parameters correlate with brain activity. This suggests that greater oscillatory dynamics are required to couple nodes and enhance inter-regional brain communication as the cognitive demands of a task increase.
2.2.2 Individual-level classification
While group-level analyses revealed significant differences between cohorts, we performed a classification analysis to test whether the inferred bifurcation parameters could serve as effective biomarkers at the individual level. A linear Support Vector Machine (SVM) classifier was trained to predict the cognitive state (seven tasks and resting state) for each subject based on their bifurcation parameter vector. The model demonstrated strong predictive power, achieving a mean accuracy of 62.71% $\pm$ 2.02% in a stratified 10-fold cross-validation scheme. This is substantially higher than the chance level of 12.50% for an 8-class problem. The performance on a held-out test set was consistent, with an accuracy of 62.63%. The confusion matrix (Fig. 3A) details the classification performance for each cohort showing that the model performed best on the Rest, Memory, and Social cohorts. The largest confusions observed align with the overlapping distributions seen in the group-level analysis (Fig. 2B).
To identify which brain regions were most critical for this classification, we conducted a permutation feature importance analysis (Fig. 3B). The results reveal a non-uniform distribution of informative regions across the brain. Notably, 10 of the top 30 regions belonged to subcortical nodes, and the only two nodes in the parcellation that belong to the Dorsal Attention network reached the top 15. By far the most influential region was the left subthalamic nucleus, followed by the right lateral occipital and the right lateral orbitofrontal cortices. In contrast, nodes from the Frontoparietal, Ventral Attention, and Default networks had few appearances on this list, suggesting their bifurcation parameters were less discriminative between the eight cohorts.
Figure 3: Individual-level classification of cohorts using bifurcation parameters. (A) Confusion matrix of the linear SVM classifier on the held-out test set for the 8-class (seven tasks and rest) classification problem. Each cell contains the raw count of predictions. The color intensity is normalized by row, representing the recall for each true class (darker green indicates higher recall). The diagonal highlights successful classification, while off-diagonal elements show specific confusion patterns between states. (B) Permutation feature importance of the top 30 most influential brain regions for the classification. The y-axis shows the mean decrease in model accuracy after permuting the values for each region 10 times. Error bars represent the standard deviation across permutations. Brain regions are colored by their corresponding brain networks as defined by the Yeo atlas (also including subcortical).
@@ 2.2.3 Brain network activations vary across tasks
To explore how different brain networks are engaged across cognitive functions, we grouped cortical nodes according to their respective brain networks, as defined by the Yeo atlas, and averaged bifurcation parameter values within each cohort. After subtracting the resting-state values from each task, we obtained the difference of mean bifurcation parameters shown in Fig. 4.
A primary finding is that nearly all differences are positive, indicating that bifurcation parameters are consistently higher during task engagement compared to rest. This suggests a widespread shift in dynamics across the cortex during active cognition.
Across tasks, we observed a remarkably consistent three-tiered hierarchy of network engagement. The top tier, showing the greatest increase in bifurcation parameters, typically comprised the Subcortical, Frontoparietal and Limbic networks. A middle tier of moderate activation consisted of the Default, Visual, and Ventral Attention networks. The bottom tier was formed by the Somatomotor network, which showed near-zero changes from rest, and the Dorsal Attention network. This well-defined hierarchy holds for all conditions except for the Social task, which uniquely disrupts the pattern. In this task, the Dorsal Attention network moves from the bottom tier to become the single most activated network (value of 0.46), showing an exceptionally large increase. The Visual network is also highly activated (0.22), showing a greater response than in any other task.
Finally, when comparing the overall magnitude of responses, the Memory task induced the smallest changes from rest across most networks. Conversely, the Emotion and Social tasks prompted the largest and most widespread increases in bifurcation parameters, suggesting the highest levels of overall network engagement.
Figure 4: Task-based differences in mean bifurcation parameters across brain networks. The mean bifurcation parameter values were calculated across cortical nodes grouped by brain network using the Yeo atlas for each task. These values were then normalized by subtracting the corresponding mean bifurcation parameter for the resting-state cohort. Each cell represents the average difference per network, with colors indicating the magnitude of the difference. Networks are organized by tiers based on overall activation intensity.
@@ 3 Discussion
We developed a novel deep-learning-based pipeline that addresses the inverse problem for an ordinary differential equation system, specifically tailored to model BOLD signals. This approach allows us to infer key dynamical parameters of a computational brain network model with remarkable efficiency and precision, relying solely on synthetic data for model training. Our results show that this model-driven inference pipeline can distinguish between different cognitive and resting states in empirical data, providing insights into the underlying brain states and supporting the use of bifurcation parameters as potential biomarkers for neural activity patterns. This capability opens a promising landscape for decoding complex task-related brain states and advancing our understanding of human neurocognition in a computationally efficient and scalable manner.
@@
Our study addresses a significant challenge in neuroscience: large volumes of training data are typically required for deep learning models to generalize effectively [Perl2020]. By generating synthetic BOLD signals derived from a simple whole-brain computational model, we circumvented the limitations associated with scarce or heterogeneous neuroimaging data, offering a controlled, high-volume dataset for training. This approach leverages synthetic data to capture essential aspects of BOLD dynamics without reliance on extensive empirical datasets, thereby avoiding potential biases from real-world data collection conditions. This methodology strengthens the potential of our model to generalize across datasets, highlighting synthetic data as a viable solution in neuroimaging applications where data scarcity often hinders model development.
@@
Previous work in computational brain network modeling has largely focused on examining the dynamics of the human brain in resting states [Deco2013, Cabral2014, Watanabe2014, Rabuffo2021]. In a foundational study, Deco et al. [Deco2017] demonstrated that the brain operates at maximum metastability during rest, characterized by continual transitions between functional network states. This metastable behavior is thought to enable optimal integration and segregation of information across regions, facilitating flexible cognitive processing. Building on this foundation, our work extends the application of the computational brain network model to task-related brain states, revealing that constructive parameters of the Hopf model —particularly the bifurcation parameters— can serve as biomarkers to differentiate distinct mental states induced by various cognitive tasks. Fig. 2B illustrates the increasing demand for inter-regional coordination and dynamic coupling as tasks deviate from the resting-state condition, accentuating the brain’s adaptive responses to varying levels of complexity. Additionally, the bifurcation parameters we obtained are predominantly small and negative, supporting the claim in this study that the brain operates near the edge of bifurcation.
@@
To the best of our knowledge, this study is among the earliest to adopt an image-based approach for processing BOLD time series using deep learning. To date, only Kancharala et al. has employed a similar strategy, using Gramian Angular Fields and Markov Transition Fields to encode BOLD signals into image representations [Kancharala2023]. That work achieved an 18% improvement in classification accuracy when distinguishing brain states induced by viewing images from standard computer vision datasets, compared to conventional time-series-based approaches. Analogously, we observed nearly a 3% decrease in normalized RMSE for bifurcation parameter prediction when using image-based transformations rather than raw time series, underscoring the advantage of this approach. By transforming temporal dynamics into spatial representations, image-based methods allow models to leverage powerful computer vision architectures adept at extracting intricate spatial features to capture complex temporal patterns across brain regions. Image-based time series methodologies in neuroimaging could pave the way for more robust and accurate tools in functional brain mapping and cognitive state classification.
@@
When analyzing the network-specific task activations shown in Fig. 4, our approach successfully links task-based activation patterns to brain networks commonly associated with these functions, aligning with previous findings in the literature [Crossley2013]:
  • The Subcortical network, a functionally diverse set of structures including the basal ganglia and thalamus, exhibited consistently high activation across all tasks. This is unsurprising given their role as central network hubs that integrate cognitive, motor, and sensory information via extensive cortico-subcortical circuits [Jansson-Boyd2024]. The results reflect their domain-general function in modulating cortical activity and supporting the transition from rest to any computationally demanding cognitive state.
  • As a core system for executive function [Niendam2012], the Frontoparietal network showed high activation across all tasks, peaking during the Emotion and Social tasks. This robust engagement underscores its essential role in integrating information and guiding behavior during demanding cognitive challenges.
  • The Limbic network, central to emotion and motivation [Themes2020], displayed a nuanced activation profile. It was highly engaged during the Motor task, a potentially surprising result suggesting a motivational or affective component to performing the cued movements. As expected, it was also strongly active during the Emotion and Gambling tasks, consistent with its role in emotional processing and reward-based decision-making.
  • Associated with internal thought [Greicius2003], the Default network showed moderate, positive activation across all tasks. This finding challenges a simple view of the network as being purely "task-negative" and supports the modern understanding that it contributes to internal cognitive processing that runs in parallel with external task demands.